Assume, unless otherwise specified, that all numbe
Answer #1

if t is the time taken to catch sammy to be caught then the distance traveled is

d= V_s t

Time

t =\frac{d}{V_s} = \frac{d}{100 (1000/3600)} = d/27.78 s .....(1)

Now the policeman catches Sammy in the same time. Police car take some time to accelerate to maximum speed and then continue on it. Distance travelled to reach the maximum speed is

V_f^2 - V_i^2 = 2 a d_1

\left ( 135 (1000/3600) \right ) - 0 = 2 (3.5) d_1

d_1 =200.9m

Time taken to reach the top speed is

t_1 = V_f / a = \frac{130 (1000/3600)}{3.5} = 10.32s

After attaining the top speed policeman has to travel certain distance to catch Sammy

d_2 = \left ( 130 (1000)/3600 \right ) t_2

t_2 = d_2 / \left ( 130 (1000)/3600 \right ) =\left ( d- 200.9 \right ) / 36.11

Total time taken is

t = t_1 + t_2 = 10. 32 + \left ( d- 200.9 \right ) / 36.11 .....(2)

Equating (1) & (2)

d/ 27.78 = 10. 32 + \left ( d- 200.9 \right ) / 36.11

Solving for d

d = 572.8 m

b)

Time taken to catch sammy is

t = 572.8/27.78 = 20.62s

c)